OpenStudy (anonymous):

The function t(x) = 3x + 5 determines how many cans of corn kernels a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew uses c(t(x)) to find the amount of money to spend on corn. The function c(x) = 2x + 3. Solve for how much money must be spent when the crew is going to work 3 shifts. 11 14 31 43

OpenStudy (anonymous):

c(t(x)), is c(x), where x = t(x)

OpenStudy (anonymous):

Write down c(x), and replace x with the equation for t(x)

OpenStudy (anonymous):

would i equal c and t together to find x ?

OpenStudy (anonymous):

c(x) = 2x+3 x = t(x) = 3x+5 2(3x+5)+3 = 6x+10+3 = 6x+13 x = 3 shifts 6(3)+10+3 = 18+10+3 = 31

OpenStudy (anonymous):

oh ok, thank you very mcuh :)

OpenStudy (anonymous):

t(3) = 3(3)+5 = 14 c(14) = 2(14)+3 = 31

OpenStudy (anonymous):

The above may be easier for you

OpenStudy (anonymous):

yes that is a lot easier thanks for helping me, i love them shortcuts